OpenStudy (anonymous):

Solve P = 4(3a - b) for a in terms of P and b. You must show all work

jimthompson5910 (jim_thompson5910):

P = 4(3a - b) P = 12a -4 b P+4b = 12a (P+4b)/12 = a a = (P+4b)/12 Answer is \[\Large a = \frac{P+4b}{12}\]